By applying the point for of OHM’s Law to a cylindrical conductor derive the expression V=IR . Assume that J and E are uniform. 


1
Expert's answer
2021-07-05T17:59:41-0400
"V=El\\J=\sigma\frac{V}{l}\\\frac{I}{A}=\sigma\frac{V}{l}\\V=I\frac{l}{\sigma A}\\V=IR"